OpenStudy (anonymous):

how to set this up? Parts for a car repair cost $175. The mechanic charges $34 per hour. If you receive an estimate for at least $226 and at most $294 for fixing the car, what is the time interval that the mechanic will be working on the job?

OpenStudy (anonymous):

remember at least and at most in the problem :)

OpenStudy (anonymous):

Well... Time*34=Price Time1=(226-175)/34 Time2=(294-175)/34 So between 1.5 to 3.5 hrs

OpenStudy (anonymous):

okay an do i have to use greater than or equal to signs?

OpenStudy (anonymous):

Sorry the equation tool isnt working for me :P I mean in mathmatical terms you could say:  Where X is the time worked on the car